A photon travels out of a material with an index of refraction 1.6 into one with an index of refraction 2.4.

(a) What is the speed of light in the first material?
m/s

(b) What is the speed of light in the second material?
m/s

(c) Say a box moving at 3.00 m/s were to slide up a ramp, and when it levels off its speed has reduced as much as the speed of the photon reduces from part (a) to part (b). How high would the ramp be?
m
Answer #1


ras speed in material v =0 1.6 V = 1.875x18 m/s to = c = 3x108 = 1.25 18 mis na 2 .u . V = 1.875*100 = v= 1-5 v₂ V2 7.25x18 n